# Quantum variational calculus on a lattice

### v2

@@ 5. Lagrangians on lattices with background metrics and gauge fields
(30) $\Delta(\phi\psi)=(\Delta\phi)\psi+\phi(\Delta\psi)+(\mathrm{id}+\sigma)(\mathrm{d}\phi+\mathrm{d}\psi)$
@@ 5.1. Scalar field theory on $X$ with a generic metric
We start by considering complex scalar field theory on an Abelian group $X$ endowed with a metric. Furthermore we choose an integration measure measure $\mu\in C(X)$ and will fix its properties as needed. In analogy to the Euclidean case, the action functional is defined as
@@
$L={1\over 2}\left(\sum_{a\in\mathcal{C}}{1\over 2}\lambda^{a}u^{2}_{a}-m^{2}u^{2}\right)\mu.$
@@
For a connected calculus, the condition in part (2) uniquely determines $\mu$, if it exists, up to an overall constant from the metric coefficients $\lambda_{a}$ and the braiding $\sigma$. This is a graph version of the way the Riemannian measure of integration is determined by the metric.
@@ 5.2. Conserved charge associated to global $U(1)$ symmetry
$j_{U(1)}=-\iota_{V}\Theta=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left(\bar{u}_{a^{-1}}u-u_{a^{-1}}\bar{u}\right)\mathrm{Vol}_{a}.$
@@
In the continuum limit, we expect $\partial_{t^{\pm 1}}\phi\mapsto\pm\partial_{0}\phi$, $\partial_{x^{\pm 1}}\phi\mapsto\pm\partial_{i}\phi$ and similar for $\phi^{*}$, that $R_{t^{-1}}\mapsto\mathrm{id}$ and $\lambda^{t^{\pm 1}}\mapsto\mathfrak{g}^{00}$, $\lambda^{x^{\pm 1}}\mapsto\mathfrak{g}^{ii}$. Therefore in both $Q_{U(1)}[\phi,\phi^{*}]$ and $J^{x}_{U(1)}[\phi,\phi^{*}]$, the second line vanishes and the first recovers the expression from Example 2.7.
@@
which as in continuum field theory correspond to 2 times the space averages in equation (22) of the real scalar field, since we now have two fields $\phi,\phi^{*}$. Again, we find similar relations as in Equations (23), (24),  (27) and  (28), but now for the energy and momentum densities directly instead of for their averages.
@@ 5.3. Complex scalar field theory on $X$ with background gauge field
Computing this expression and assuming the same relation between the measure $\mu$, inverse metric $\lambda^{a}$ and braiding $\sigma$ as in Proposition 5.1, we recover exactly the operator in (34), so that the latter becomes the covariantized Klein-Gordon equation $(\Delta_{\alpha}+m^{2})\phi=0$. We have done this for $X$ an Abelian group but as with the end of Section 5.1, there is also a version on any graph.

### v3

@@ 5. Lagrangians on lattices with background metrics and gauge fields
(28) $\Delta(\phi\psi)=(\Delta\phi)\psi+\phi(\Delta\psi)-\frac{1}{2}(\mathrm{id}+\sigma)(\mathrm{d}\phi+\mathrm{d}\psi)$
@@ 5.1. Scalar field theory on $X$ with a generic metric
We start by considering complex scalar field theory on an Abelian group $X$ endowed with a generic metric. Furthermore we choose an integration measure measure $\mu\in C(X)$ and will fix its properties as needed. In analogy to the Euclidean case, the action functional is defined as
@@
$L={\mu\over 2}\left(\sum_{a\in\mathcal{C}}{1\over 2}\lambda^{a}u^{2}_{a}-m^{2}u^{2}\right).$
@@
For a connected calculus, the condition in part (2) uniquely determines $\mu$, if it exists, up to an overall constant from the metric coefficients $\lambda_{a}$ and the braiding $\sigma$. This is a graph version of the way the Riemannian measure of integration is determined by the metric in Riemannian geometry.
@@ 5.2. Conserved charge associated to global $U(1)$ symmetry
$j_{U(1)}=-\iota_{V}\Theta=\frac{{\mathfrak{i}}\varphi}{4}\sum_{a\in\mathcal{C}}R_{a^{-1}}\left(\mu\lambda^{a}\right)\left(\bar{u}_{a^{-1}}u-u_{a^{-1}}\bar{u}\right)\mathrm{Vol}_{a},$
@@
In the continuum limit, we expect $\partial_{t^{\pm 1}}\phi\mapsto\pm\partial_{0}\phi$, $\partial_{x^{\pm 1}}\phi\mapsto\pm\partial_{i}\phi$ and similar for $\phi^{*}$, that $R_{t^{-1}}\mapsto\mathrm{id}$ and $\lambda^{t^{\pm 1}}\mapsto\mathfrak{g}^{00}$, $\lambda^{x^{\pm 1}}\mapsto\mathfrak{g}^{ii}$. Therefore in both $Q_{U(1)}[\phi,\phi^{*}]$ and $J^{x}_{U(1)}[\phi,\phi^{*}]$, the second line vanishes and the first recovers the expression from Example 2.7 for $g$ the Euclidean or Minkowski metrics.
@@
with now $A\in\mathbb{C}$, which as in continuum field theory correspond to 2 times the spatial averages in equation (21) of the real scalar field, since we now have two fields $\phi,\phi^{*}$. Again, we find similar relations as in equations (22), (23),  (25) and  (26), but now for the energy and momentum densities directly instead of for their averages.
@@ 5.3. Complex scalar field theory on $X$ with background gauge field
Computing this expression and assuming the same relation between the measure $\mu$, inverse metric $\lambda^{a}$ and braiding $\sigma$ as in Proposition 5.1, we recover exactly the operator in (32), so that the latter becomes the covariantised Klein-Gordon equation $(\Delta_{\alpha}+m^{2})\phi=0$. We have done this for $X$ an Abelian group but as with the end of Section 5.1, there is also a version on any graph.